OpenStudy (errinss):

A flag pole is 12 feet tall. Its shadow is 10 feet long. What is the approximate distance from the top of the flag pole to the tip of the pole's shadow? A. 22 feet B. 16 feet C. 11 feet D. 7 feet

OpenStudy (errinss):

@Cheez @rishavraj

OpenStudy (welshfella):

use the pythagoras theorem x^2 = 10^2 + 12^2 find x.

rishavraj (rishavraj):

same method :))

OpenStudy (errinss):

\[12^2+10^2=15.62^2\]

OpenStudy (errinss):

or in this case i would round up to 16^2

rishavraj (rishavraj):

so the answer is 16 :)

OpenStudy (errinss):

awesome